# Fantastic Weights and How to Find Them: Where to Prune in Dynamic Sparse Training

### v1

@@ Abstract
Dynamic Sparse Training (DST) is a rapidly evolving area of research that seeks to optimize the sparse initialization of a neural network by adapting its topology during training. It has been shown that under specific conditions, DST is able to outperform dense models. The key components of this framework are the pruning and growing criteria, which are repeatedly applied during the training process to adjust the network’s sparse connectivity. While the growing criterion’s impact on DST performance is relatively well studied, the influence of the pruning criterion remains overlooked. To address this issue, we design and perform an extensive empirical analysis of various pruning criteria to better understand their effect on the dynamics of DST solutions. Surprisingly, we find that most of the studied methods yield similar results. The differences become more significant in the low-density regime, where the best performance is predominantly given by the simplest technique: magnitude-based pruning. The code is provided at https://github.com/alooow/fantastic_weights_paper
@@ 1 Introduction
Modern deep learning solutions have demonstrated exceptional results in many different disciplines of science brown2020language ; fawzi2022discovering ; jumper2021highly . However, they come at the cost of using an enormous number of parameters. Consequently, compression methods aim to significantly reduce the model size without introducing any loss in the performance han2015learning ; cheng2017survey .
One approach to sparsifying neural networks is pruning, in which a portion of the weights is removed at the end of the training based on some predefined importance criterion lecun1989optimal ; hassibi1993optimal ; han2015learning . More recently, frankle2018lottery have found that iterative pruning joined with cautions parameter re-initialization can identify sparse subnetworks that are able to achieve similar performance to their dense counterparts when trained from scratch. This result, known as the lottery ticket hypothesis, has launched subsequent research into methods for identifying and training models that are sparse already at initialization lee2019snip ; wang2020picking ; zhou2019deconstructing ; ramanujan2020s ; sreenivasan2022rare .
@@
Motivated by the above, we take a closer look at the current DST techniques. Most research in this domain is devoted to investigating the best growth criterion as the most influential factor in the performance of DST bellec2018deep ; evci2020rigging ; dettmers2019sparse ; atashgahi2022brain , disregarding any potential contribution coming from the choice of the weight removal algorithm. In this study, we address this issue by taking a complementary approach and focusing on the pruning criteria instead. A pruning criterion in DST serves as a measure of weight importance and hence is a proxy of the “usefulness" of a particular connection. Note that the importance of a connection in DST can differ from standard post-training pruning, as the role of weights can change throughout training due to the network’s plasticity and adaptation. Weights deemed unimportant in one step can become influential in later phases of the training.
@@
The insights from our research identify that the simplest magnitude pruning is still the optimal choice, despite a large amount of alternatives present in the literature. This drives the community’s attention to carefully examine any new adaptation criteria for Dynamic Sparse Training.
@@ 2 Related Work
Pruning and Sparse Training. A common way of reducing the neural network size is pruning, which removes parameters or entire blocks of layers from the network. In its classical form, pruning has been extensively studied in the context of compressing post-training models janowsky1989pruning ; mozer1989using ; lecun1989optimal ; hassibi1993optimal ; han2015learning ; narang2017exploring ; li2016pruning ; molchanov2019importance ; gale2019state – see e.g. hoefler2021sparsity ; blalock2020state for an overview and survey. Interestingly, mocanu2016topological demonstrated that a sparse neural network can match and even outperform its corresponding dense neural network equivalent if its sparse connectivity is designed in a sensitive manner. Recently, a similar result was obtained by the lottery ticket hypothesis and follow-up research, which showed that there exist sparse subnetworks that can be trained in isolation to the same performance as the dense networks frankle2018lottery ; zhou2019deconstructing ; wu2019sparsemask . In an effort to find these subnetworks without the need for dense training, different techniques have been proposed over the years lee2019snip ; wang2020picking ; tanaka2020pruning , including random selection ramanujan2020s ; liu2022unreasonable . Such approaches are commonly referred to as Sparse Training, or Static Sparse Training to emphasize that the sparse structure stays the same throughout the training. Dynamic Sparse Training. In DST, the neural network structure is constantly evolving by pruning and growing back weights during training mocanu2018scalable ; bellec2018deep ; dettmers2019sparse ; evci2020rigging ; yuan2021mest ; wortsman2019discovering ; atashgahi2022brain ; dai2019nest ; kundu2023autosparse . The key motivation behind DST is not only to provide compression but also to increase the effectiveness and robustness of the deep learning models without the need for overparametrization liu2021we . The capability of DST has recently been an area of interest. In evci2022gradient , authors indicate that DST is able to outperform static sparse training by assuring better gradient flow. DST has also been reported to achieve high performance in Reinforcement Learning graesser2022state ; sokar2021dynamic ; grooten2023automatic and Continual Learning sokar2021spacenet , with ongoing research into applicability in NLP liu2023sparsity . Some attention has also been given to the topological properties of the connectivity patterns produced by DST. In particular, liu2020topological , investigated the structural features of DST. However, they focus only on one method, the Sparse Evolutionary Training (SET) procedure mocanu2018scalable .
@@ 3.1 Dynamic Sparse Training
Figure 1: The test accuracy versus density computed on the studied models for different pruning criteria (due to space limits we only present here six plots, see Figure 9 in Appendix C for the remaining two setups). The first row represents the results obtained by random growth, while the second corresponds to gradient growth. Note the logarithmic scale in the x-axis. The performance of the dense model is indicated by the horizontal dashed line. In almost every case all pruning criteria perform well, regardless of the chosen model and growth criterion, at the same time surpassing the static initialization (in light blue).
@@ 3.2 Pruning Criteria
Given a pruning fraction $\rho_{t}$, the pruning criterion $\mathcal{C}$ determines the importance score $s(\theta^{l}_{i,j})$ for each weight and prunes the ones with the lowest score. We use local pruning, which means that the criterion is applied to each layer separately  [^1] We adapt this setup as it is commonly used in DST approaches, and has a reduced risk of entirely disconnecting the network – see Appendix H for a more detailed discussion.. For brevity, we will slightly abuse the notation and write $\theta$ instead of $\theta^{l}_{i,j}$ from now on. Below, we discuss the most commonly used pruning criteria in DST that are the subject of the conducted study.
@@ 4 Methodology
The common sense expectation would be that the more elaborate pruning criteria, which combine the weight magnitude with the gradient information, will perform better than the techniques that rely only on the weight’s absolute value. Since the gradient may provide information on how the weight will change in the future, it might be more suitable in the DST approach, where connectivity is constantly evolving. This is especially promising considering the effectiveness of the gradient-based approach in estimating the importance in the growth criterion.
@@
In the second research question, we focus on one of the most important design choices: the topology update period $\Delta t$. The setting of this update period determines how often the network structure is adapted during training. Within a fixed dataset and batch size, using a relatively low value means that the topology is changed very frequently in comparison to the optimization update of the weights. This poses a potential risk of not letting the newly added connections reach their full potential. On the other hand, a high value gives the weights enough opportunity to become important but may not allow the sparse network to adjust its structure based on the data characteristics. Finding a balance between this exploitation-exploration conflict is an imperative task in any DST framework.
(a) all densities
(b) density $\leq 0.2$
(c) density $>0.2$
@@
Finally, we take a deep dive into the structures produced by the various pruning criteria. Please note that while the previous questions, Q1 and Q2, may explain how the pruning criterion choice affects performance and update frequency in DST, they do not indicate whether the mask solutions obtained by those criteria are genuinely different. In order to assess the diversity of the studied methods in terms of the produced sparse connectivity structures, we need to compare the sets of weights selected for pruning by each criterion under the same network state. In addition, we also investigate the similarity between the final masks obtained at the end of training and how close they are to their corresponding sparse initializations. This allows us to assess whether the DST topology updates are indeed meaningful in changing the mask structure. For both cases, we incorporate a common score of the proximity of sets, known as the Jaccard index (or Intersection-Over-Union). We compute it separately for each layer and average the result:
@@ 5.1 Setup of the Experiments
We perform our analysis using eight different models, including small- and large-scale MLPs and Convolutional Nets. The small-MLP is a 4-layer network with hidden size of at most $256$ (trained on the tabular Higgs dataset baldi2014searching ). The large-MLP also consists of 4 layers (latent dimension size up to $1024$ neurons) and is evaluated on the CIFAR10 krizhevsky2009learning . The convolutional architectures are: a small 3-layer CNN (CIFAR10), LeNet5-Caffe lecun1998gradient (FashionMNIST xiao2017fashion ), ResNet56 he2016deep (CIFAR10, CIFAR100), VGG-16 simonyan2014very (CIFAR100) and EfficientNet tan2019efficientnet (Tiny-ImageNet le2015tiny ). In addition, on a selected density, we also consider the ResNet50 model on ImageNet imagenet . We summarize all the architectures in Appendix A.
@@
We train the models using the DST framework on a set of predefined model densities. We use adjustment step $\Delta t=800$ and initial pruning fraction $\rho=0.5$ with cosine decay. As the growth criterion, we investigate both the random and gradient-based approaches, as those are the most common choices in the literature. In all cases, we ensure that within each model, the training hyperparameters setup for every criterion is the same. Each performance is computed by averaging 5 runs, except for ImageNet, for which we use 3 runs. The ResNet-50 model is trained using density $0.2$ and gradient growth (we select the gradient-based growth method as it is known to provide good performance in this setup evci2020rigging ). Altogether we performed approximately 7000 runs of dense, static sparse, and DST models, which in total use 8 (9 including ImageNet) different dataset-model configurations, and jointly took approximately 5275 hours on GPU. [^3] For training details please refer to Appendix A.2.
5.2 Impact on the Performance of DST
Firstly, we observe that the differences between the pruning criteria are usually most distinctive in low densities. Furthermore, in that setting, the dynamic sparse training framework almost always outperforms the static sparse training approach. This is important, as in DST research, high sparsity regime is of key interest. Surprisingly, the more elaborate criteria using gradient-based approaches usually either perform worse than the simple magnitude score, or do not hold a clear advantage over it. For instance, the Taylor-based $\mathcal{C}_{\text{SNIP}}$ criterion, despite being still better than the static initialization, consistently achieves results similar or worse than $\mathcal{C}_{\text{Magnitude}}$. This is especially visible in the DST experiments using gradient growth (note ResNet56, MLP-CIFAR10, and EfficientNet). The only exception is the VGG-16 network. In addition, we also observe that $\mathcal{C}_{\text{SNIP}}$ gives better performance when used in a fine-tuning setup for EfficientNet - see Appendix G. Consequently, this Taylor-approximation-based criterion, although being well suited for standard post-training pruning molchanov2019importance and selecting sparse initialization masks lee2019snip seems not to be the optimal choice in dynamic sparse training approaches on high sparsity. The $\mathcal{C}_{\text{RSensitivity}}$, on the other hand, can outperform other methods but is better suited for high densities. Indeed, it is even the best choice for densities larger than $0.15$ for large-MLP on CIFAR10 and ResNet models. Finally, the $\mathcal{C}_{\text{MEST}}$ criterion also is not superior to the $\mathcal{C}_{\text{Magnitude}}$, usually leading to similar results. [^4] Note that we also study the impact of the $\lambda$ hyperparameter of the $\mathcal{C}_{\text{MEST}}$ criterion in Appendix A.3. In general, we see that $\lambda\rightarrow 0$ leads to results equal to $\mathcal{C}_{\text{Magnitude}}$, while $\lambda\rightarrow\infty$ prioritizes only the gradient magnitude and degrades the performance. We do not see any clear improvement for values lying between those two extremes. We also present the results obtained for ResNet-50 on ImageNet with density $0.2$ in Figure 2. We observe that, again, the $\mathcal{C}_{\text{SNIP}}$ criterion leads to the worst performance. The $\mathcal{C}_{\text{RSensitivity}}$, $\mathcal{C}_{\text{SET}}$, and $\mathcal{C}_{\text{MEST}}$ methods achieve high results but not clearly better than the $\mathcal{C}_{\text{Magnitude}}$ criterion.
Figure 5: Left: The loss gap (training loss -- validation loss) over time for different pruning criteria, the static sparse model, the dense model, and the dense model with dropout p=0.05 on the large-MLP. Right: The test accuracy obtained for each criterion. The DST methods exhibit lower gaps than the dense model. At the same time they achieve better results than the dense and dense+dropout models.
Within each studied density, growing criterion, and model, we rank the pruning criteria and then calculate their average ranks. To rigorously establish the statistical significance of our findings, we compute the Critical Distance Diagram for those ranks, using the Nemenyi post-hoc test demvsar2006statistical with p-value $0.05$ and present it in Figure 3. We observe that in the low-density regime (Figure 2(b)), the $\mathcal{C}_{\text{Magnitude}}$ criterion achieves the best performance, as given by the lowest average rank. Furthermore, in such a case, we also note that other predominately magnitude-based criteria, such as $\mathcal{C}_{\text{SET}}$ and $\mathcal{C}_{\text{MEST}}$ are not significantly different than $\mathcal{C}_{\text{Magnitude}}$, while the remaining gradient-based approaches are clearly worse. This confirms the overall effectiveness of the weight’s magnitude as an importance score in sparse regimes. [^5] In Appendix F we loosely hypothesize that the low performance of gradient-based approaches may be due to the high variance in the gradient. Interestingly, when larger densities are involved, the best choice is given by $\mathcal{C}_{\text{RSensitivity}}$, while the performance of $\mathcal{C}_{\text{magnitude}}$ deteriorates. This may suggest that the information from the gradient is more reliable in such a case. The gradient-based methods seem to perform better in denser setups.
Finally, we notice that in certain situations, the DST framework can lead to better results than the original dense model (see small-MLP, or ResNet56 on CIFAR100 in Figure 1). We argue that this is due to a mix of a regularization effect and increased expressiveness introduced by removing and adding new connections in the network – see Figure 5. Moreover, we also observe that within the setup studied in Figure 1, there is almost no difference between choosing the random or gradient growth criterion. The second one achieves slightly better accuracy for the convolutional models (see e.g., density $0.2$ on ResNet-56, as well as improved stability for EfficientNet). The similarity between those two growth criteria and the surprising effectiveness of DST in comparison to dense models have also been previously observed in the context of Reinforcement Learning graesser2022state .
5.3 Sensitivity to Update Period
In this section, we analyze the sensitivity of each pruning criterion to the update period $\Delta t$. This value describes how often the adjustments are made in training and hence can affect the weights’ importance scores. We fix the density to $0.2$ and $0.1$ and the initial pruning fraction $\rho$ to $0.5$. We choose the densities $0.2$ and $0.1$ as they correspond to rather high sparsity levels and generally provide reasonable performance. Next, we investigate different update periods on the models using CIFAR10 dataset. [^6] The update period is only meaningful when considered together with the size of the training set. Therefore it is important to evaluate the models on the same dataset. We start from $\Delta t=25$ and increase it by a factor of $2$ up to $\Delta t=6400$.
@@ 5.3 Sensitivity to Update Period
The results are presented in Figure 4. For the MLP model, we clearly see the behavior described in Section 4: if the update period $\Delta t$ is too small, the performance will suffer. The best update period setting seems to be $\Delta t=800$ for the random growth, regardless of the pruning method. For gradient growth, this value also gives a good performance. However, larger pruning periods (e.g. $\Delta t=1600$) are even better. Note that we use a batch size of $128$, hence for $\Delta t=800$ the connectivity is changed approximately once per 2 epochs. [^7] Since $800\cdot 128=102,400$ samples, which is just over twice the training set size of CIFAR10 ($50,000$). Similarly, less frequent updates are also beneficial in the ResNet56 model, although there is not as much consensus between the different pruning criteria. Even performing just one topology update every $\Delta t=6400$ iterations (approximately once per 16 epochs) still gives excellent performance. At the same time, not performing any topology updates at all (i.e., static sparse training) deteriorates performance, as shown by the cyan dashed line in Figure 4. Furthermore, we observe that the gradient-based pruning criteria seem to suffer much more than the magnitude criteria when paired with a small update period (consider the plots for $\Delta t=25$), especially when combined with the gradient growth. [^8] In Appendix D we also study the exploration imposed by the pruning criteria using the ITOP ratio liu2021we . This suggests that it is safer not to use those techniques together or be more rigorous about finding the right update period in such a case.
@@
Figure 6: (a) The mean normalized Jaccard index between the sets of weights chosen for removal during the first update of the sparse connectivity, computed for the MLP, CNN, and ResNet-56 models on CIFAR10. The $J_{r}$ indicates the expected overlap of random subsets and serves as a reference point. We estimate this value by computing the mean pair-wise Jaccard index of the random pruning criterion with different sparse initializations. (b) The same index computed between the masks obtained by different pruning criteria at the end of training. The rows and columns represent the pruning criteria between which the index is computed. The $J_{init}$ denotes the expected overlap of random masks and serves as a reference point. We estimate this value by computing the mean pair-wise Jaccard index of the different sparse initializations.

### v2

@@ Abstract
Dynamic Sparse Training (DST) is a rapidly evolving area of research that seeks to optimize the sparse initialization of a neural network by adapting its topology during training. It has been shown that under specific conditions, DST is able to outperform dense models. The key components of this framework are the pruning and growing criteria, which are repeatedly applied during the training process to adjust the network’s sparse connectivity. While the growing criterion’s impact on DST performance is relatively well studied, the influence of the pruning criterion remains overlooked. To address this issue, we design and perform an extensive empirical analysis of various pruning criteria to better understand their impact on the dynamics of DST solutions. Surprisingly, we find that most of the studied methods yield similar results. The differences become more significant in the low-density regime, where the best performance is predominantly given by the simplest technique: magnitude-based pruning.
@@ 1 Introduction
Modern deep learning solutions have demonstrated exceptional results in many different disciplines of science brown2020language ; fawzi2022discovering ; jumper2021highly . However, they come at the cost of using an enormous number of parameters. Consequently, compression methods aim to significantly reduce the model size without introducing any loss in performance han2015learning ; cheng2017survey .
One approach to sparsifying neural networks is pruning, in which a portion of the weights is removed at the end of the training based on some predefined importance criterion lecun1989optimal ; hassibi1993optimal ; han2015learning . More recently, frankle2018lottery have found that iterative pruning joined with cautious parameter re-initialization can identify sparse subnetworks that are able to achieve similar performance to their dense counterparts when trained from scratch. This result, known as the lottery ticket hypothesis, has launched subsequent research into methods for identifying and training models that are sparse already at initialization lee2019snip ; wang2020picking ; zhou2019deconstructing ; ramanujan2020s ; sreenivasan2022rare .
@@
Motivated by the above, we take a closer look at the current DST techniques. Most of the research concerning design choices in DST focuses on the analysis of methods where variations occur in the growth criterion, while the pruning criterion becomes intertwined with other design considerations bellec2018deep ; evci2020rigging ; dettmers2019sparse ; atashgahi2022brain . In this study, we address this issue by taking a complementary approach and focusing on the pruning criteria instead. A pruning criterion in DST serves as a measure of weight importance and hence is a proxy of the "usefulness" of a particular connection. Note that the importance of a connection in DST can differ from standard post-training pruning, as the role of weights can change throughout training due to the network’s plasticity and adaptation. Weights deemed unimportant in one step can become influential in later phases of the training.
@@
The insights from our research identify that the simplest magnitude pruning is still the optimal choice, despite a large amount of alternatives present in the literature. This drives the community’s attention to carefully examine any new adaptation criteria for Dynamic Sparse Training. [^1] The code is provided at https://github.com/alooow/fantastic_weights_paper.
@@ 2 Related Work
Pruning and Sparse Training. A common way of reducing the neural network size is pruning, which removes parameters or entire blocks of layers from the network. In its classical form, pruning has been extensively studied in the context of compressing post-training models janowsky1989pruning ; mozer1989using ; lecun1989optimal ; hassibi1993optimal ; han2015learning ; narang2017exploring ; li2016pruning ; molchanov2019importance ; gale2019state – see e.g. hoefler2021sparsity ; blalock2020state for an overview and survey. Interestingly, mocanu2016topological demonstrated for the first time in the literature that a sparse neural network can match and even outperform its corresponding dense neural network equivalent if its sparse connectivity is designed in a sensitive manner. Recently, a similar result was obtained by the lottery ticket hypothesis and follow-up research, which showed that there exist sparse subnetworks that can be trained in isolation to the same performance as the dense networks frankle2018lottery ; zhou2019deconstructing ; wu2019sparsemask . In an effort to find these subnetworks without the need for dense training, different techniques have been proposed over the years lee2019snip ; wang2020picking ; tanaka2020pruning , including random selection ramanujan2020s ; liu2022unreasonable . Such approaches are commonly referred to as Sparse Training, or Static Sparse Training to emphasize that the sparse structure stays the same throughout the training. Interestingly, frankle2020pruning find out that the sparse initializations produced by the static approaches are invariant to parameter reshuffling and reinitialization within a layer and that even pre-training magnitude pruning performs quite well in such a setup.
Dynamic Sparse Training. In DST, the neural network structure is constantly evolving by pruning and growing back weights during training mocanu2018scalable ; bellec2018deep ; dettmers2019sparse ; evci2020rigging ; yuan2021mest ; wortsman2019discovering ; atashgahi2022brain ; dai2019nest ; kundu2023autosparse . The key motivation behind DST is not only to provide compression but also to increase the effectiveness and robustness of the deep learning models without the need for overparametrization liu2021we . The capability of DST has recently been an area of interest. In evci2022gradient , authors indicate that DST is able to outperform static sparse training by assuring better gradient flow. DST has also been reported to achieve high performance in Reinforcement Learning graesser2022state ; sokar2021dynamic ; grooten2023automatic and Continual Learning sokar2021spacenet , with ongoing research into applicability in NLP liu2023sparsity . Some attention has also been given to the topological properties of the connectivity patterns produced by DST. In particular, liu2020topological , investigated the structural features of DST. However, they focus only on one method, the Sparse Evolutionary Training (SET) procedure mocanu2018scalable .
@@ 3.2 Pruning Criteria
Given a pruning fraction $\rho_{t}$, the pruning criterion $\mathcal{C}$ determines the importance score $s(\theta^{l}_{i,j})$ for each weight and prunes the ones with the lowest score. We use local pruning, which means that the criterion is applied to each layer separately. [^2] We adapt this setup as it is commonly used in DST approaches, and has a reduced risk of entirely disconnecting the network – see Appendix I for a more detailed discussion.. For brevity, we will slightly abuse the notation and write $\theta$ instead of $\theta^{l}_{i,j}$ from now on. Below, we discuss the most commonly used pruning criteria in DST that are the subject of the conducted study.
@@ 4 Methodology
Figure 1: The test accuracy versus density computed on the studied models for different pruning criteria (due to space limits we only present here six plots, see Figure 10 in Appendix D for the remaining two setups). The first row represents the results obtained by random growth, while the second corresponds to gradient growth. Note the logarithmic scale in the x-axis. The performance of the dense model is indicated by the horizontal dashed line. In almost every case all pruning criteria perform well, regardless of the chosen model and growth criterion. At the same time, they surpass the static initialization (in light blue).
The common sense expectation would be that the more elaborate pruning criteria, which combine the weight magnitude with the gradient information, will perform better than the techniques that rely only on the weight’s absolute value. Since the gradient may provide information on how the weight will change in the future, it indicates the trend the weight might follow in the next optimization updates. Therefore it may be more suitable in the DST approach, where connectivity is constantly evolving, making insights into future updates potentially beneficial. This is especially promising considering the effectiveness of the gradient-based approach in estimating the importance in the growth criterion.
@@
In the second research question, we focus on one of the most important design choices: the topology update period $\Delta t$. The setting of this update period determines how often the network structure is adapted during training. When considered together with the dataset– and batch-size, this hyperparameter can be interpreted as the number of data samples the model trains on before performing the mask update. Therefore, within a fixed dataset and batch size, using a relatively low value means that the topology is changed very frequently in comparison to the optimization update of the weights. This poses a potential risk of not letting the newly added connections reach their full potential. On the other hand, a high value gives the weights enough opportunity to become important but may not allow the sparse network to adjust its structure based on the data characteristics. Finding a balance between this exploitation-exploration conflict is an imperative task in any DST framework.
@@
Finally, we take a deep dive into the structures produced by the various pruning criteria. Please note that while the previous questions, Q1 and Q2, may explain how the pruning criterion choice affects performance and update frequency in DST, they do not indicate whether the mask solutions obtained by those criteria are significantly different. In order to assess the diversity of the studied methods in terms of the produced sparse connectivity structures, we need to compare the sets of weights selected for pruning by each criterion under the same network state. In addition, we also investigate the similarity between the final masks obtained at the end of training and how close they are to their corresponding sparse initializations. This allows us to assess whether the DST topology updates are indeed meaningful in changing the mask structure. For both cases, we incorporate a common score of the proximity of sets, known as the Jaccard index (or Intersection-Over-Union). We compute it separately for each layer and average the result:
@@ 5.1 Setup of the Experiments
We perform our analysis using eight different models, including small- and large-scale MLPs and Convolutional Nets. The small-MLP is a 4-layer network with hidden size of at most $256$, (trained on the tabular Higgs dataset baldi2014searching ). The large-MLP also consists of 4 layers (latent dimension size up to $1024$ neurons) and is evaluated on the CIFAR10 krizhevsky2009learning . The convolutional architectures are: a small 3-layer CNN (CIFAR10), LeNet5-Caffe lecun1998gradient (FashionMNIST xiao2017fashion ), ResNet56 he2016deep (CIFAR10, CIFAR100), VGG-16 simonyan2014very (CIFAR100) and EfficientNet tan2019efficientnet (Tiny-ImageNet le2015tiny ). In addition, on a selected density, we also consider the ResNet50 model on ImageNet imagenet . We summarize all the architectures in Appendix A. While our primary focus is on vision and tabular data, we also evaluate the pruning criteria on the fine-tuning task of ROBERTa Large liu2019roberta using the CommonsenseQA datasettalmor2018commonsenseqa adapted from the Sparsity May Cry (SMC) Benchmark liu2023sparsity . In contrast to the setup studied throughout this work, this experiment involves a fine-tuning type of problem. In consequence, we discuss it in Appendix C and focus on training models from scratch for the remainder of the paper.
We train the models using the DST framework on a set of predefined model densities. We use adjustment step $\Delta t=800$ and initial pruning fraction $\rho=0.5$ with cosine decay. As the growth criterion, we investigate both the random and gradient-based approaches, as those are the most common choices in the literature. Following yuan2021mest , we reuse the gradient computed during the backward pass when calculating the scores in the gradient-based pruning criteria. In all cases, we ensure that within each model, the training hyperparameters setup for every criterion is the same. Each performance is computed by averaging 5 runs, except for ImageNet, for which we use 3 runs. The ResNet50 model is trained using density $0.2$ and gradient growth (we select the gradient-based growth method since is known to provide good performance in this task evci2020rigging ). Altogether we performed approximately 7000 runs of dense, static sparse, and DST models, which in total use 8 (9 including ImageNet) different dataset-model configurations, and jointly took approximately 5275 hours on GPU. [^4] For training details please refer to Appendix A.2.
5.2 Impact on the Performance of DST
(a) all densities
(b) density $\leq 0.2$
(c) density $>0.2$
Firstly, we observe that the differences between the pruning criteria are usually most distinctive in low densities. Furthermore, in that setting, the dynamic sparse training framework almost always outperforms the static sparse training approach. This is important, as in DST research, high sparsity regime is of key interest. Surprisingly, the more elaborate criteria using gradient-based approaches usually either perform worse than the simple magnitude score, or do not hold a clear advantage over it. For instance, the Taylor-based $\mathcal{C}_{\text{SNIP}}$ criterion, despite being still better than the static initialization, consistently achieves results similar or worse than $\mathcal{C}_{\text{Magnitude}}$. This is especially visible in the DST experiments using gradient growth (note ResNet56, MLP-CIFAR10, and EfficientNet). The only exception is the VGG-16 network. In addition, we also observe that $\mathcal{C}_{\text{SNIP}}$ gives better performance when used in a fine-tuning setup for EfficientNet - see Appendix H. Consequently, this Taylor-approximation-based criterion, although being well suited for standard post-training pruning molchanov2019importance and selecting sparse initialization masks lee2019snip seems not to be the optimal choice in dynamic sparse training approaches on high sparsity. The $\mathcal{C}_{\text{RSensitivity}}$, on the other hand, can outperform other methods but is better suited for high densities. Indeed, it is even the best choice for densities larger than $0.15$ for large-MLP on CIFAR10 and ResNet models. Finally, the $\mathcal{C}_{\text{MEST}}$ criterion also is not superior to the $\mathcal{C}_{\text{Magnitude}}$, usually leading to similar results. [^5] We study the impact of the $\lambda$ hyperparameter of the $\mathcal{C}_{\text{MEST}}$ criterion in Appendix A.3. In general, we see that $\lambda\rightarrow 0$ leads to results equal to $\mathcal{C}_{\text{Magnitude}}$, while $\lambda\rightarrow\infty$ prioritizes only the gradient magnitude and degrades the performance. We do not see any clear improvement for values lying between those two extremes. We also present the results obtained for ResNet-50 on ImageNet with density $0.2$ in Figure 2. We observe that, again, the $\mathcal{C}_{\text{SNIP}}$ criterion leads to the worst performance. The $\mathcal{C}_{\text{RSensitivity}}$, $\mathcal{C}_{\text{SET}}$, and $\mathcal{C}_{\text{MEST}}$ methods achieve high results but not clearly better than the $\mathcal{C}_{\text{Magnitude}}$ criterion.
Figure 4: Left: The loss gap (validation loss -- training loss) over time for different pruning criteria, the static sparse model, the dense model, and the dense model with dropout $p=0.05$ on the large-MLP. Right: The test accuracy obtained for each criterion.
Within each studied density, growing criterion, and model, we rank the pruning criteria and then calculate their average ranks. To rigorously establish the statistical significance of our findings, we compute the Critical Distance Diagram for those ranks, using the Nemenyi post-hoc test demvsar2006statistical with p-value $0.05$ and present it in Figure 3. We observe that in the low-density regime (Figure 2(b)), the $\mathcal{C}_{\text{Magnitude}}$ criterion achieves the best performance, as given by the lowest average rank. Furthermore, in such a case, we also note that other predominately magnitude-based criteria, such as $\mathcal{C}_{\text{SET}}$ and $\mathcal{C}_{\text{MEST}}$ are not significantly different than $\mathcal{C}_{\text{Magnitude}}$, while the remaining gradient-based approaches are clearly worse. This confirms the overall effectiveness of the weight’s magnitude as an importance score in sparse regimes. See also Appendix G for the discussion of the effect of batch size on the obtained results. Interestingly, when larger densities are involved, the best choice is given by $\mathcal{C}_{\text{RSensitivity}}$. Additionally, the average rank of $\mathcal{C}_{\text{magnitude}}$ deteriorates. This may suggest that the information from the gradient is more reliable in such a case. The gradient-based methods seem to perform better in denser setups.
Finally, we notice that in certain situations, the DST framework can lead to better results than the original dense model (see small-MLP, or ResNet56 on CIFAR100 in Figure 1). We argue that this is due to a mix of a regularization effect and increased expressiveness introduced by removing and adding new connections in the network – see Figure 4. Moreover, we also observe that within the setup studied in Figure 1, there is almost no difference between choosing the random or gradient growth criterion. The second one achieves slightly better accuracy for the convolutional models (see, e.g., density $0.2$ on ResNet-56, as well as improved stability for EfficientNet). The similarity between those two growth criteria and the surprising effectiveness of DST in comparison to dense models have also been previously observed in the context of Reinforcement Learning graesser2022state .
5.3 Sensitivity to Update Period
In this section, we analyze the sensitivity of each pruning criterion to the update period $\Delta t$. This value describes how often the adjustments are made in training and hence can affect the weights’ importance scores. We fix the density to $0.2$ and $0.1$ and the initial pruning fraction $\rho$ to $0.5$. We choose the densities $0.2$ and $0.1$ as they correspond to rather high sparsity levels and generally provide reasonable performance. Next, we investigate different update periods on the models using CIFAR10 dataset. [^6] The update period is only meaningful when considered together with the size of the training set. Therefore it is important to evaluate models of different sizes on the same dataset. We start from $\Delta t=25$ and increase it by a factor of $2$ up to $\Delta t=6400$.
@@ 5.3 Sensitivity to Update Period
The results are presented in Figure 5. For the MLP model, we clearly see the behavior described in Section 4: if the update period $\Delta t$ is too small, the performance will suffer. The best update period setting seems to be $\Delta t=800$ for the random growth, regardless of the pruning method. For gradient growth, this value also gives a good performance. However, larger pruning periods (e.g. $\Delta t=1600$) are even better. [^7] See also Appendix J for an analysis of the effect of batch size and pruning schedule on these results. Note that we use a batch size of $128$, hence for $\Delta t=800$ the connectivity is changed approximately once per 2 epochs. [^8] Since $800\cdot 128=102,400$ samples, which is just over twice the training set size of CIFAR10 ($50,000$). Similarly, less frequent updates are also beneficial in the ResNet56 model, although there is not as much consensus between the different pruning criteria. Even performing just one topology update every $\Delta t=6400$ iterations (approximately once per 16 epochs) still gives excellent performance. At the same time, not performing any topology updates at all (i.e., static sparse training) deteriorates performance, as shown by the cyan dashed line in Figure 5. Furthermore, we observe that the gradient-based pruning criteria seem to suffer much more than the magnitude criteria when paired with a small update period (consider the plots for $\Delta t=25$), especially when combined with the gradient growth. [^9] In Appendix E we also study the exploration imposed by the pruning criteria using the ITOP ratio liu2021we . This suggests that it is safer not to use those techniques together or be more rigorous about finding the right update period in such a case.
Figure 6: (a) The mean normalized Jaccard index between the sets of weights chosen for removal during the first update of the sparse connectivity, computed for the MLP, CNN, and ResNet-56 models on CIFAR10. The $J_{r}$ indicates the expected overlap of random subsets and serves as a reference point. We estimate this value by computing the mean pair-wise Jaccard index of the random pruning criterion with different sparse initialization. (b) The same index computed between the masks obtained by different pruning criteria at the end of training. The rows and columns represent the pruning criteria between which the index is computed. The $J_{init}$ denotes the expected overlap of random masks and serves as a reference point. We estimate this value by computing the mean pair-wise Jaccard index of the different sparse initializations.